OpenStudy (anonymous):

Which function is equivalent to y = 16^0.5x? The options are... 1. y=0.5^16x 2. y=8x 3. y=4^2x 4. y=4x Please help me ;/

OpenStudy (anonymous):

16^0.5 =4

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

you just solve it? fml x_x

OpenStudy (anonymous):

4^2 = 16 so it is 4^2(0.5x) = 4^x

OpenStudy (anonymous):

Is the X in the original problem in the exponent?